(Calculating free cash flows​) At​ present, Solartech Skateboards is considering expanding its product line to include​ gas-powered skateboards;​ however, it is questionable how well they will be received by skateboarders. Although you feel there is a70percent chance you will sell 8,000 of these per year for 10 years​ (after which time this project is expected to shut down because​ solar-powered skateboards will become more​ popular), you also recognize that there is a 15 percent chance that you will only sell 5,000 and also a 15 percent chance you will sell 15,000.

The gas skateboards would sell for $100 each and have a variable cost of $50 each. Regardless of how many you​ sell, the annual fixed costs associated with production would be $140,000. In​ addition, there would be an initial expenditure of ​$1,100,000 associated with the purchase of new production equipment which will be depreciated using the bonus depreciation method in year 1. Because of the number of stores that will need​ inventory, the working capital requirements are the same regardless of the level of sales. This project will require a​ one-time initial investment of $30,000 in net working​ capital, and​ working-capital investment will be recovered when the project is shut down.​ Finally, assume that the​ firm's marginal tax rate is 22 percent.

a. What is the initial outlay associated with the​ project?

b. What are the annual free cash flows associated with the project for years​ 1, and 2 through 9 under each sales​ forecast? What are the expected annual free cash flows for year​ 1, and years 2 through​ 9?

c. What is the terminal cash flow in year 10​ (that is, what is the free cash flow in year 10 plus any additional cash flows associated with the termination of the​ project)?

d. Using the expected free cash​ flows, what is the​ project's NPV given a required rate of return of 12 ​percent? What would the​ project's NPV be if 8,000 skateboards were​ sold?
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Required Return =12%=0.12
a Initial Outlay
A Production Equipment $1,100,000
B Investment in Net working Capital $30,000
C=A+B Initial Outlay Associated with the project $1,130,000
b Sales Forecast Sales Forecast Sales Forecast
D Sales Forecast                       8,000                      5,000                    15,000
Cash Flow in Year 1:
E=D*($100-$50) Contribution Margin $400,000 $250,000 $750,000
F Fixed Costs -$140,000 -$140,000 -$140,000
G Depreciation -$1,100,000 -$1,100,000 -$1,100,000
H=E+F+G Operating Income Before tax -$840,000 -$990,000 -$490,000
I=H*(1-0.22) After Tax Operating Income -$655,200 -$772,200 -$382,200
J Add: Depreciation (Non cash expense) $1,100,000 $1,100,000 $1,100,000
K=I+J Operating Cash Flow $444,800 $327,800 $717,800
Probability of Cash Flow                          0.70                        0.15                        0.15 SUM
Cash Flow*Probability $311,360 $49,170 $107,670 $468,200
Expected Cash Flow in Year1 $468,200
Cash Flow in Year 2-9
D Sales Forecast                       8,000                      5,000                    15,000
E=D*($100-$50) Contribution Margin $400,000 $250,000 $750,000
F Fixed Costs -$140,000 -$140,000 -$140,000
H=E+F Operating Income Before tax $260,000 $110,000 $610,000
I=H*(1-0.22) After Tax Operating Income $202,800 $85,800 $475,800
K Operating Cash Flow $202,800 $85,800 $475,800
Probability of cash flow                          0.70                        0.15                        0.15 SUM
Cash Flow*Probability $141,960 $12,870 $71,370 $226,200
Expected Cash Flow in Year2-9 $226,200
c Terminal Cash Flow in Year10
K Operating Cash Flow $202,800 $85,800 $475,800
L Release of net working capital $30,000 $30,000 $30,000
M Terminal Cash Flow in Year10 $232,800 $115,800 $505,800
Probability of Cash Flow                          0.70                        0.15                        0.15 SUM
Cash Flow*Probability $162,960 $17,370 $75,870 $256,200
Expected Cash Flow in Year10 $256,200
8,000 $400,000 $140,000 $260,000 $202,800 $202,800 0.70 $141,960 $226,200 5,000 $250.00 $140,000 $110,000 $85,800 $85,800 0.1